How many liters of a 40% acid solution and a 55% acid solution must be mixed to produce 60 liters of a 45% acid solution?
Write 2 equations and solve using substitution or elimination?
Not out of a text book!

Answer by checkley71(8403)

.4X+.55(60-X)=60*.45 OR .4X+33-.55X=27 OR -.15X=-6 OR X=6/.15 OR X=40 LITERS OF 40% SOLUTION THEN 60-40=20 LITERS OF 55% SOLUTION ARE NEEDED TO MAKE 60 LITERS OF A 45% MIXTURE.